OpenStudy (anonymous):

Prove. (1-tanx)(1-cotx)=2-secxcscx

OpenStudy (anonymous):

(1-tanx)(1-cotx)=1-cotx-tanx+1=2-(cotx+tanx)=2-((cosx/sinx)+(sinx/cosx))=2-(1/sinxcosx)=2-secxcscx

OpenStudy (anonymous):

so you factored first?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Distributive_property \[(1-\tan x)(1-\cot x)=1-\cot x-\tan x+1=2-(\cot x+\tan x)=2-[(\frac{ \cos x }{ \sin x })+(\frac{ \sin x }{ \cos x })]=2-(\frac{ \cos ^{2}x +\sin ^{2}x}{ sinxcosx })=2-(cscx)(secx)\]